mhchen:

Show that if p(x) is a polynomial with a nonzero constant term, and if A is a square matrix for which p(A) = 0, then A is invertible.

Narad:

The polynomial is \[p(x)= a _{n}x ^{n}+a _{n-1}x ^{n-1}+.....+a_{1}x+a _{0}\] Then, \[p(A)= a _{n}A ^{n}+a _{n-1}A ^{n-1}+......a _{1}A+a _{0}I=0\] so, \[a _{n}A ^{n}+a _{n-1}A ^{n-1}+........a _{1}A=-a _{0}I\] \[A(a _{n}A ^{n-1}+ a _{n-1}A ^{n-2}+....a _{1}I)=-a _{0}I\] If A is invertible \[AA ^{-1}=I\] Hence, you can calculate \[A ^{-1}\]

mhchen:

@Narad okay that makes a it easier to see. Thanks

Narad:

You are welcome

mhchen:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Narad The polynomial is \[p(x)= a _{n}x ^{n}+a _{n-1}x ^{n-1}+.....+a_{1}x+a _{0} = 0\] Then, \[p(A)= a _{n}A ^{n}+a _{n-1}A ^{n-1}+......a _{1}A+a _{0}I=0\] \[a _{n}A ^{n}+a _{n-1}A ^{n-1}+........a _{1}A=-a _{0}I\] \[\frac{a _{n}}{-a_{0}}A ^{n}+ \frac{a _{n-1}}{-a_{0}}A ^{n-1}+....\frac{a _{1}}{-a_{0}}A=I\] \[A(\frac{a _{n}}{-a_{0}}A ^{n-1}+ \frac{a _{n-2}}{-a_{0}}A ^{n-1}+....\frac{a _{1}}{-a_{0}}I)=I\] If A is invertible \[AA ^{-1}=I\] Hence, you can calculate \[A ^{-1}\] \(\color{#0cbb34}{\text{End of Quote}}\)